OpenStudy (anonymous):

Differentiate the given function: f(x)= ln (x+1)/(x-1) I get -2/(x-1)^2

OpenStudy (anonymous):

-2/ (x^2-1)

OpenStudy (anonymous):

I used the quotient rule, but not sure what to do with the ln

OpenStudy (anonymous):

did u have? ln(x+1/x-1)?

OpenStudy (anonymous):

if yes, then d/dx (lnx)= 1/x ... easy

OpenStudy (anonymous):

the ln is before the fraction

OpenStudy (anonymous):

Quotient rule u= ln (x+1) v= ln (x-1)

OpenStudy (anonymous):

If ln is over the whole thing ln [(x+1)/(x-1)] rewrite as [ln (x+1)]/[ln (x-1)]

OpenStudy (anonymous):

I don't know if this will help ln x+1 x-1 x+1/x-1 one big parenthesis around this fraction

OpenStudy (anonymous):

Can be rewritten as \[\ln (x+1)\div \ln (x-1)\]

OpenStudy (anonymous):

Top is u, bottom v

OpenStudy (anonymous):

I have not done it that way

OpenStudy (anonymous):

the u and v

OpenStudy (anonymous):

what shorthand you use for quotient rule f(x) or what?

OpenStudy (anonymous):

ok for quotient rule, I can do that but not sure what to do with the ln. After the quotient rule I get (x-1)-(x+1) / (x-1)^2

OpenStudy (anonymous):

Excuse me for using u and v. Let u=ln (x+1), let v=ln (x-1) u'=?, v'=? u' means u prime or derivative of u The derivative of the whole thing is \[(u'v-uv')\div(v ^{2})\]

OpenStudy (anonymous):

I am sorry but I am having a real hard time understanding this math. I have a book that has this problem worked out. I just don't understand why certain things are done the way they are. The last section we went over was dervatives and I did really well with them. These I am not.

OpenStudy (anonymous):

OK, so write one or two steps from book and say what you don't understand

OpenStudy (anonymous):

1st step f'(x) = 1/(x+1/x-1) d/dx (x+1/x-1) why did they put 1 over the problem.

OpenStudy (anonymous):

next step x-1/x+1 [(x-1)(1) - (x+1)(1)//(x-1)^2 the 2nd half I get, that is the quotient rule

OpenStudy (anonymous):

they are treating the whole fraction thingy as a single number. Let (x+1)/(x-1) be u (sorry) f' of ln u=1/u

OpenStudy (anonymous):

Deal with one question at a time or everything would get confused

OpenStudy (anonymous):

ok with step one 1/(x+1/x-1) then I guess they took the bottom # and flipped and multiplied. Is that how the problem got switched for step 2

OpenStudy (anonymous):

OK, back to original question what might make it a little confusing is what is called chain rule. the f' (x) = 1/[x+1/(x-1)] du in addition to this you must find the derivative of the inner function (x+1)/(x-1). Find the derivative of that and that is your du. This is a complicated function. I can explain more after you look at it.

OpenStudy (anonymous):

why can't I use the quotient rule

OpenStudy (anonymous):

by putting the 1 / .... is that taking the ln out of the problem

OpenStudy (anonymous):

There is a lot going on in this problem. It is not that you can't use quotient rule. In fact they are using quotient rule. But there are some intricacies that make their answer difficult to understand. Start from the top. Ask one piece by one piece what you don't understand.

OpenStudy (anonymous):

ok 1st step 1/ (x+1/x-1) if I use the quotient rule derv. is x+1/x-1

OpenStudy (anonymous):

never mind that isn't correct.

OpenStudy (anonymous):

Just write the book line by line and I will do the play by play

OpenStudy (anonymous):

1st step f'(x) = 1/(x+1/x-1) d/dx (x+1/x-1) 2nd step x-1/x+1 [(x-1)(1) - (x+1)(1)/(x-1)^2 step 3 -2/(x+1)(x-1) That is all steps

OpenStudy (anonymous):

Are you familiar with the thing called the chain rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There approach is good. Now that I see it. They considered the ln (that thing) ln (that thing)=[1/(that thing)][multiplied by derivative of that thing] They didn't find it necessary to do the quotient rule to find f'(x), but in finding the derivative of (that thing) they did the quotient rule. Derivative of ln doesn't require work, you just put 1/something

OpenStudy (anonymous):

ok I half way understand that but how did they get x-1/x+1 in step 2 before the quotient rule

OpenStudy (anonymous):

sometimes when they work the problems out in the book it doesn't explain what they are doing. If I understand the work it does not confuss me.

OpenStudy (anonymous):

1/(B/C)=C/B

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You have fractional property at the bottom is flipped.

OpenStudy (anonymous):

ok you flip it to get rid of the 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok on step 2 x-1/x+1 [(x-1)(1)-(x+1)(1)/(x-1)^2 ok I can see the quotient rule what are they doine with the x-1/x+1, are they multipling that by the quotient?

OpenStudy (anonymous):

never mind something just clicked, I see it now. They cross multiplied